In a garden, 2/3 of the flowers are roses. Of the roses in the garden, 5/12 are yellow and the rest are red. What fraction of the flowers are red roses?
2/3 16/24, 24/24-16/24=8/24=1/3
5/12of the 1/3= are yellow roses
7/12 of the remainder 1/3 are red
7/12*1/3=7/36 are red roses

Answer by richwmiller(17219)

In a garden, 2/3 of the flowers are roses. Of the roses in the garden, 5/12 are yellow and the rest are red. What fraction of the flowers are red roses?
f= flowers
o= roses
y= yellow roses
r= red roses
2f/3=o
5o/12=y
7o/12=r
2f/3=o
7(2f/3)/12=r
r = (7 f)/18
7/18 of the flowers are red roses